Determine the comparable interest rate for a $70,000 loan when the quoted information is 11.2% + 0.5 pt + $200. (Round your answer to two decimal places.)
Answer by rfer(16322)

70000*.112=7840
70000*0.005=350
7840+350+200=8390
8390/70000=0.1198=0.12=11.98%
I can't tell if they want just the decimal or the %
so I gave you both.